# Influence of initiators on the tipping point in the extended Watts model

### v1

@@ I Introduction
In recent years, there has been a growing interest among researchers in social contagion processes in networks, such as the spread of opinions and the adoption of innovations Jalili and Perc (2017); Guilbeault et al. (2018). In a pioneering study, Watts proposed a social contagion model (hereafter referred to as the Watts model) on a network, where each node representing an individual can be either active, indicating that it has adopted an innovation, or inactive Watts (2002); Watts and Dodds (2007). An inactive node changes its state to active by social influence if the fraction of active nodes among its neighbors exceeds a predetermined adoption threshold. The adoption threshold of a node represents its sensitivity to social influence: a lower adoption threshold signifies that fewer active neighbors are required for the node to become active, whereas a higher adoption threshold signifies that more active neighbors are required for the node to become active. In networks, the activation of a very small fraction of nodes can cause large cascades, where almost all nodes eventually become active if the mean adoption threshold is below a certain value Watts (2002). The transition from a small cascade region, where cascades hardly spread, to a large cascade region, where cascades spread extensively, is often associated with a jump in the expected size of cascades. Following the initial study by Watts Watts (2002), the relationship between the likelihood of large cascades and the structure of complex networks has been extensively investigated Ikeda et al. (2010); Hackett et al. (2011); Hackett and Gleeson (2013); Gleeson (2008); Payne et al. (2009); Dodds and Payne (2009); Payne et al. (2011); Galstyan and Cohen (2007); Nematzadeh et al. (2014); Curato and Lillo (2016); Yağan and Gligor (2012); Zhuang et al. (2017); Unicomb et al. (2019). For example, studies on the Watts model in clustered networks Ikeda et al. (2010); Hackett et al. (2011); Hackett and Gleeson (2013) and correlated networks Gleeson (2008); Payne et al. (2009); Dodds and Payne (2009); Payne et al. (2011) have demonstrated that highly clustered structures and positively degree-correlated structures, respectively, facilitate the induction of large cascades. Furthermore, some studies have proposed variants of the Watts model to model more complex contagion processes Dodds and Watts (2004, 2005); Melnik et al. (2013); Hurd and Gleeson (2013); Ruan et al. (2015); Kobayashi (2015); Karsai et al. (2016); Lee and Holme (2017); Huang et al. (2016); Oh and Porter (2018); Chung et al. (2019).
@@
In this paper, we investigate the effect of the influence of seeds on the tipping point in cascade dynamics using the extended Watts model Nishioka and Hasegawa (2022). In this extended model, nodes with connections to seeds (referred to as seed neighbors) have a different adoption threshold than ordinary nodes. We apply the tree approximation to the extended Watts model on Erdős–Rényi random graphs (ERRGs) to determine the tipping points for different combinations of adoption thresholds for seed neighbors and ordinary nodes. We observe that the occurrence of a tipping point and the scale of cascades are dependent on two factors: whether seed neighbors form a giant component (when the seed fraction is large enough to allow cascades of activations between seed neighbors), and whether ordinary nodes form a giant component (when activated nodes trigger further activations of ordinary nodes). For example, if seed neighbors form a giant component, a large cascade can occur between them, even if ordinary nodes have such a high adoption threshold that almost none of them is ever activated. In the case where a giant component of seed neighbors and that of ordinary nodes coexist, two tipping points can appear. Our tree approximation demonstrates that for the extended Watts model with a gap between the adoption thresholds of seed neighbors and ordinary nodes, the active node fraction increases discontinuously at two different tipping points as the seed fraction increases. Monte Carlo simulations clearly show that the first cascade, occurring at a small tipping point, occurs in the giant component of seed neighbors, while the second cascade, occurring at a larger tipping point, extends into the giant component of ordinary nodes.
@@ II Model
This study employs the extended Watts model introduced in Nishioka and Hasegawa (2022). Let us consider a network in which each node is in one of two states: active or inactive. A fraction $\rho_{0}$ of nodes are randomly selected as seeds, which are initially active. Among other nodes, which are initially inactive, seed neighbors are nodes that are connected to one or more seeds, while ordinary nodes are nodes that are not connected to any seeds  [^1] In Ref. Nishioka and Hasegawa (2022), seed neighbors were called direct neighbors, while ordinary nodes were called indirect neighbors. However, these terms were confusing and have been corrected in this paper.. The cascade dynamics begins from an initial state in which seeds are active and non-seed nodes are inactive. A seed neighbor of degree $k$ changes its state to active if the proportion of $k$ adjacent nodes that are active exceeds a predetermined threshold $\theta_{\rm SN}$, that is, $m\geq k\theta_{\rm SN}$, where $m$ denotes the number of active neighbors. Similarly, an ordinary node of degree $k$ changes its state to active if the proportion of $k$ adjacent nodes that are active exceeds a predetermined threshold $\theta_{\rm ON}$, that is, $m\geq k\theta_{\rm ON}$. Once activated, nodes always remain active, and the updating process continues until no additional nodes become active. The fraction of active nodes in the final state is denoted by $\rho_{\infty}(\geq\rho_{0})$, where $\rho_{\infty}$ describes the scale of the cascade: $\rho_{\infty}\approx\rho_{0}$ when cascades triggered by seeds hardly spread, and $\rho_{\infty}\approx 1$ when cascades triggered by seeds cause almost all nodes in the network to be active in the final state.
@@ III Results
First, we briefly review the case of $\theta_{\rm SN}=\theta_{\rm ON}=\theta$, where the extended Watts model corresponds to the original Watts model with threshold $\theta$. Figure 1(a) plots the active node fraction $\rho_{\infty}$ as a function of the seed fraction $\rho_{0}$ for the Watts model on an ERRG with average degree $\langle k\rangle=6$. For $\theta=0.55$ and $0.65$, the active node fraction increases continuously with $\rho_{0}(<\rho_{0,{\rm c}})$, jumping to $\rho_{\infty}\approx 1$ at the tipping point $\rho_{0}=\rho_{0,{\rm c}}(\theta)$. A larger adoption threshold $\theta$ results in a tipping point $\rho_{0,{\rm c}}(\theta)$ appearing at a larger seed fraction. For large values of $\theta$, no discontinuous transitions are observed, as seen in the case of $\theta=0.75$. Figure 1(b) displays the fixed points $r_{\infty}$ of Eq. (5) as a function of $\rho_{0}$. When $\theta$ is small, there are three fixed points for small $\rho_{0}$ and one fixed point for large $\rho_{0}$, signifying that a discontinuous transition occurs at some seed fraction. In contrast, when $\theta$ is large, there is always one fixed point regardless of $\rho_{0}$, signifying that there is no jump in $\rho_{\infty}$. In general, the seed fraction required for activations to spread increases with an increase in $\theta$. However, cascades of activations are possible only on the connected components of non-seed nodes, which become smaller and split as the number of seeds increases. The disappearance of jumps in $\rho_{\infty}$ for large values of $\theta$ is attributed to the fact that the connected components of non-seed nodes are split into finite ones due to an excessive number of seeds before activations begin to spread. For a large cascade to occur in the network, non-seed nodes must be sufficiently connected to form a giant component, occupying a finite fraction of nodes in the network, when the seed fraction becomes large enough to allow activations to propagate.
@@
Figure 2: Active node fraction $\rho_{\infty}$ as a function of $\rho_{0}$ for the extended Watts model when the adoption threshold of ordinary nodes is set to (a) $\theta_{\rm ON}=5/10$ and (b) $\theta_{\rm ON}=9/10$. We employed an ERRG with $\langle k\rangle=10$ and evaluated $\rho_{\infty}$ using Eqs. (4) and (5). The red dotted line, green solid line, and blue dashed line represent the results when the adoption threshold of seed neighbors is $\theta_{\rm SN}=0.2,0.4$, and $0.6$, respectively. The symbols represent the simulation results for ERRGs with $10^{5}$ nodes, averaged over $10^{3}$ samples ($10^{2}$ graphs $\times$ $10$ initial states). The theoretical lines are in good agreement with the simulation results.
Next, we consider the case of $\theta_{\rm SN}\neq\theta_{\rm ON}$. Figures 2(a) and (b) plot the active node fraction $\rho_{\infty}$ as a function of the seed fraction $\rho_{0}$ for the extended Watts model on an ERRG with average degree $\langle k\rangle=10$, where the adoption threshold of ordinary nodes is set to $\theta_{\rm ON}=5/10$ and $9/10$, respectively. As illustrated in Fig. 2(a), when $\theta_{\rm ON}$ is small (i.e., ordinary nodes can easily become active), then when seed neighbors are activated by seeds, further activations are triggered by active seed neighbors, resulting in an activation cascade of ordinary nodes. In this case, the active node fraction jumps to $\rho_{\infty}\approx 1$ at the tipping point $\rho_{0,{\rm c}}$. In contrast, as illustrated in Fig. 2(b), when $\theta_{\rm ON}$ is high (i.e., ordinary nodes do not easily become active), activation cascades may not occur between ordinary nodes even when most seed neighbors become active. When $\theta_{\rm SN}=0.2$ and $0.4$, $\rho_{\infty}$ increases discontinuously at the tipping point but does not reach $\rho_{\infty}\approx 1$, and increases continuously as $\rho_{0}$ increases above the tipping point. At the tipping point, ordinary nodes are rarely activated, and activations spread only over the giant component formed by seed neighbors.
@@
In the present model, it is natural to assume that the scale of cascades is related to whether nodes of each type form a giant component at different seed fractions. We refer to the connected components of seeds, the connected components of seed neighbors, the connected components of ordinary nodes, and the connected components of seed neighbors and ordinary nodes as S-components, SN-components, ON-components, and (SN+ON)-components, respectively. Using generating functions, we derive the size of these components and their associated critical points (Appendix A). Figure 3(a) presents the normalized giant S-component size $S_{\rm S}$, giant SN-component size $S_{\rm SN}$, giant ON-component size $S_{\rm ON}$, and giant (SN+ON)-component size $S_{\rm SN+ON}$ of the ERRG with $\langle k\rangle=10$ as a function of the seed fraction $\rho_{0}$. The giant S-component appears above a critical point $\rho_{0,{\rm c}}^{\rm S}=0.1$ when $\rho_{0}$ increases ($S_{\rm S}=0$ for $\rho_{0}\leq\rho_{0,{\rm c}}^{\rm S}$ and $S_{\rm S}>0$ for $\rho_{0}>\rho_{0,{\rm c}}^{\rm S}$). The giant ON-component and the giant (SN+ON)-component occupy most of the network when the seed fraction is small; however, their sizes decrease with $\rho_{0}$ and disappear at $\rho_{0,{\rm c}}^{\rm ON}\approx 0.207$ and $\rho_{0,{\rm c}}^{\rm SN+ON}=0.9$, respectively ($S_{\rm ON}=0$ for $\rho_{0}\geq\rho_{0,{\rm c}}^{\rm ON}$ and $S_{\rm SN+ON}=0$ for $\rho_{0}\geq\rho_{0,{\rm c}}^{\rm SN+ON}$). When $\rho_{0,{\rm c}}^{\rm ON}<\rho_{0}<\rho_{0,{\rm c}}^{\rm SN+ON}$, where $S_{\rm ON}=0$ but $S_{\rm SN+ON}>0$, the giant (SN+ON)-component is mostly composed of seed neighbors.
@@
The giant SN-component behaves non-monotonically with an increasing seed fraction $\rho_{0}$. As $\rho_{0}$ increases from zero, the number of seed neighbors first increases, and the giant SN-component is formed at a smaller critical point ($\rho_{0,{\rm c}}^{\rm SN}\approx 0.011$). However, as the seed fraction continues to increase, the connectivity between seed neighbors is obstructed by seeds. As the seed fraction exceeds a larger critical point ($\rho_{0,{\rm c}}^{\rm SN}\approx 0.900$), the giant SN-component disappears. Thus, the giant SN-component exists only in the region where $\rho_{0}$ is between the two values of $\rho_{0,{\rm c}}^{\rm SN}$  ($S_{\rm SN}>0$ if $0.011\lessapprox\rho_{0}\lessapprox 0.900$ and $S_{\rm SN}=0$ otherwise). Figure 3(b) plots the critical point for each type of giant component in the ($\rho_{0}$, $\langle k\rangle$) plane. The figure indicates that when the average degree of the ERRG is $\langle k\rangle>\langle k\rangle_{\rm c}\approx 2.580$, each type of giant component behaves qualitatively the same as in Fig. 3(a). However, when $1<\langle k\rangle<\langle k\rangle_{\rm c}$, seed neighbors never form a giant component. In the absence of a giant SN-component, a jump in $\rho_{\infty}$ does not occur unless there is a cascade of activations between ordinary nodes. However, when the giant SN-component is present, it is possible for a cascade to occur among seed neighbors, resulting in a jump in $\rho_{\infty}$ even if ordinary nodes are rarely activated.
@@
Figure 4: (a) Active node fraction $\rho_{\infty}$ and (b) fixed points $r_{\infty}$ of Eq. (5) for the extended Watts model on an ERRG with $\langle k\rangle=10$. The adoption threshold of seed neighbors is set to $\theta_{\rm SN}=1/4$. In (a), the red dotted line, black solid line, and blue dashed line represent $\rho_{\infty}$ when $\theta_{\rm ON}=1/5$, $2/3$, and $9/10$, respectively. In (b), the red open squares, black filled circles, and blue open triangles represent $r_{\infty}$ when $\theta_{\rm ON}=1/5$, $2/3$, and $9/10$, respectively. The open circles in (a) represent the simulation results for ERRGs with $10^{5}$ nodes, averaged over $10^{3}$ samples ($10^{2}$ graphs $\times$ $10$ initial states). The simulation results are in good agreement with the theoretical line, exhibiting two jumps of $\rho_{\infty}$.
@@
Figure 5: Size distribution of active components (black squares), active SN-components (red circles), and active ON-components (blue triangles) for (a) $\rho_{0}=0.03<\rho_{0,{\rm c}}^{\rm 1st}$ (below the first tipping point), (b) $\rho_{0,{\rm c}}^{\rm 1st}<\rho_{0}=0.06<\rho_{0,{\rm c}}^{\rm 2nd}$ (above the first tipping point and below the second tipping point), and (c) $\rho_{0}=0.09>\rho_{0,{\rm c}}^{\rm 2nd}$ (above the second tipping point). The adoption thresholds of seed neighbors and ordinary nodes are set to $\theta_{\rm SN}=1/4$ and $\theta_{\rm ON}=2/3$, respectively. Each distribution, defined as the number of corresponding components of size $s$ per node, is obtained from Monte Carlo simulations of $10^{6}$ samples ($10^{4}$ initial states $\times$ $10^{2}$ graphs) for the extended Watts model on ERRGs with $10^{5}$ nodes.
The coexistence of a giant SN-component and giant ON-component suggests that a large cascade on each giant component can occur at different tipping points. Here, we demonstrate that for the extended Watts model, a tipping point exhibiting a jump in $\rho_{\infty}$ can appear twice. Let us consider the extended Watts model on an ERRG with $\langle k\rangle=10$, where a giant SN-component and giant ON-component coexist for $0.011\lessapprox\rho_{0}\lessapprox 0.207$. Figure 4(a) illustrates the active node fraction $\rho_{\infty}$ when $\theta_{\rm SN}=1/4$ and $\theta_{\rm ON}=1/5,2/3$, and $9/10$. When $\theta_{\rm ON}=2/3$ (i.e., there is a gap between the adoption thresholds of seed neighbors and ordinary nodes), there are five fixed points for $0.043\lessapprox\rho_{0}\lessapprox 0.047$, three fixed points for $0.047\lessapprox\rho_{0}\lessapprox 0.066$, and one fixed point for $\rho_{0}\gtrapprox 0.066$ (Fig. 4(b)). Consequently, the active node fraction exhibits a jump at $\rho_{0}=\rho_{0,{\rm c}}^{\rm 1st}\approx 0.047$ and at $\rho_{0}=\rho_{0,{\rm c}}^{\rm 2nd}\approx 0.066$ as $\rho_{0}$ increases.

### v2

@@ I Introduction
In recent years, there has been a growing interest among researchers in social contagion processes in networks, such as the spread of opinions and the adoption of innovations Jalili and Perc (2017); Guilbeault et al. (2018). In a pioneering study, Watts proposed a social contagion model (hereafter referred to as the Watts model) on a network, where each node representing an individual can be either active, indicating that it has adopted an innovation, or inactive Watts (2002); Watts and Dodds (2007). An inactive node changes its state to active by social influence if the fraction of active nodes among its neighbors exceeds a predetermined adoption threshold. The adoption threshold of a node represents its sensitivity to social influence: a lower adoption threshold signifies that fewer active neighbors are required for the node to become active, whereas a higher adoption threshold signifies that more active neighbors are required for the node to become active. In networks, the activation of a very small fraction of nodes can cause large cascades, where almost all nodes eventually become active if the mean adoption threshold is below a certain value Watts (2002). The transition from a small cascade region, where cascades hardly spread, to a large cascade region, where cascades spread extensively, is often associated with a jump in the expected size of cascades. Following the initial study by Watts Watts (2002), the relationship between the likelihood of large cascades and the structure of complex networks has been extensively investigated Ikeda et al. (2010); Hackett et al. (2011); Hackett and Gleeson (2013); Gleeson (2008); Payne et al. (2009); Dodds and Payne (2009); Payne et al. (2011); Galstyan and Cohen (2007); Nematzadeh et al. (2014); Curato and Lillo (2016); Yağan and Gligor (2012); Zhuang et al. (2017); Unicomb et al. (2019). For example, studies on the Watts model in clustered networks Ikeda et al. (2010); Hackett et al. (2011); Hackett and Gleeson (2013) and correlated networks Gleeson (2008); Payne et al. (2009); Dodds and Payne (2009); Payne et al. (2011) have demonstrated that highly clustered structures and positively degree-correlated structures, respectively, facilitate the induction of large cascades. Furthermore, some studies have proposed variants of the Watts model to model more complex contagion processes Dodds and Watts (2004, 2005); Melnik et al. (2013); Hurd and Gleeson (2013); Ruan et al. (2015); Kobayashi (2015); Karsai et al. (2016); Lee and Holme (2017); Huang et al. (2016); Wang et al. (2017); Oh and Porter (2018); Chung et al. (2019).
@@
In this paper, we investigate the effect of the influence of seeds on the tipping point in cascade dynamics using the extended Watts model Nishioka and Hasegawa (2022). In this extended model, nodes with connections to seeds (referred to as seed neighbors) have a different adoption threshold than ordinary nodes. We apply the tree approximation to the extended Watts model on Erdős–Rényi random graphs (ERRGs) to determine the tipping points for different combinations of adoption thresholds for seed neighbors and ordinary nodes. We observe that the occurrence of a tipping point and the scale of cascades are dependent on two factors: whether seed neighbors form a giant component (when the seed fraction is large enough to allow cascades of activations between seed neighbors), and whether ordinary nodes form a giant component (when activated nodes trigger further activations of ordinary nodes). For example, if seed neighbors form a giant component, a large cascade can occur between them, even if ordinary nodes have such a high adoption threshold that almost none of them is ever activated. In the case where a giant component of seed neighbors and that of ordinary nodes coexist, two tipping points can appear. Our tree approximation demonstrates that for the extended Watts model with a gap between the adoption thresholds of seed neighbors and ordinary nodes, the active node fraction increases discontinuously at two different tipping points as the seed fraction increases. Monte Carlo simulations clearly show that the first cascade, occurring at a smaller tipping point, takes place in the giant component of seed neighbors, while the second cascade, occurring at a larger tipping point, extends into the giant component of ordinary nodes.
@@ II Model
This study employs the extended Watts model introduced in Nishioka and Hasegawa (2022). Let us consider a network in which each node is in one of two states: active or inactive. A fraction $\rho_{0}$ of nodes are randomly selected as seeds, which are initially active. Among other nodes, which are initially inactive, seed neighbors are nodes that are connected to one or more seeds, while ordinary nodes are nodes that are not connected to any seeds  [^1] In Ref. Nishioka and Hasegawa (2022), seed neighbors were called direct neighbors, while ordinary nodes were called indirect neighbors. However, these terms were confusing and have been corrected in this paper.. Whether seed neighbors and ordinary nodes become active depends on the predetermined adoption thresholds, $\theta_{\rm SN}$ and $\theta_{\rm ON}$, respectively. A seed neighbor of degree $k$ changes its state to active if the proportion of $k$ adjacent nodes that are active exceeds a predetermined threshold $\theta_{\rm SN}$, that is, $m\geq k\theta_{\rm SN}$, where $m$ denotes the number of active neighbors. Similarly, an ordinary node of degree $k$ changes its state to active if the proportion of $k$ adjacent nodes that are active exceeds a predetermined threshold $\theta_{\rm ON}$, that is, $m\geq k\theta_{\rm ON}$.
We distinguish between the adoption thresholds of seed neighbors and ordinary nodes to assume $\theta_{\rm SN}<\theta_{\rm ON}$. Seed neighbors can be considered as individuals who are directly influenced by the initial information source, which may be the first to adopt the innovation with great interest as technology enthusiasts or visionaries. Their proximity to the source makes them more likely to be easily activated, reflecting their lower adoption threshold $\theta_{\rm SN}$. Ordinary nodes, on the other hand, represent individuals who are not directly connected to the initial information source. These nodes typically require a larger amount of influence to become activated due to their distance from the source, reflecting their higher adoption threshold $\theta_{\rm ON}$.
The cascade dynamics begins from an initial state in which seeds are active and non-seed nodes are inactive. Once activated, nodes always remain active, and the updating process continues until no additional nodes become active. The fraction of active nodes in the final state is denoted by $\rho_{\infty}(\geq\rho_{0})$, where $\rho_{\infty}$ describes the scale of the cascade: $\rho_{\infty}\approx\rho_{0}$ when cascades triggered by seeds hardly spread, and $\rho_{\infty}\approx 1$ when cascades triggered by seeds cause almost all nodes in the network to be active in the final state.
@@ III Results
First, we briefly review the case of $\theta_{\rm SN}=\theta_{\rm ON}=\theta$, where the extended Watts model corresponds to the original Watts model with threshold $\theta$. Figure 1(a) plots the active node fraction $\rho_{\infty}$ as a function of the seed fraction $\rho_{0}$ for the Watts model on an ERRG with average degree $\langle k\rangle=6$. For $\theta=0.55$ and $0.65$, the active node fraction increases continuously with $\rho_{0}(<\rho_{0,{\rm c}})$, jumping to $\rho_{\infty}\approx 1$ at the tipping point $\rho_{0}=\rho_{0,{\rm c}}(\theta)$. A higher adoption threshold $\theta$ results in a tipping point $\rho_{0,{\rm c}}(\theta)$ appearing at a larger seed fraction. For large values of $\theta$, no discontinuous transitions are observed, as seen in the case of $\theta=0.75$. Figure 1(b) displays the fixed points $r_{\infty}$ of Eq. (5) as a function of $\rho_{0}$. When $\theta$ is low, there are three fixed points for small $\rho_{0}$ and one fixed point for large $\rho_{0}$, signifying that a discontinuous transition occurs at some seed fraction. In contrast, when $\theta$ is high, there is always one fixed point regardless of $\rho_{0}$, signifying that there is no jump in $\rho_{\infty}$. In general, the seed fraction required for activations to spread increases with an increase in $\theta$. However, cascades of activations are possible only on the connected components of non-seed nodes, which become smaller and split as the number of seeds increases. The disappearance of jumps in $\rho_{\infty}$ for large values of $\theta$ is attributed to the fact that the connected components of non-seed nodes are split into finite ones due to an excessive number of seeds before activations begin to spread. For a large cascade to occur in the network, non-seed nodes must be sufficiently connected to form a giant component, occupying a finite fraction of nodes in the network, when the seed fraction becomes large enough to allow activations to propagate.
@@
Figure 2: Active node fraction $\rho_{\infty}$ as a function of $\rho_{0}$ for the extended Watts model when the adoption threshold of ordinary nodes is set to (a) $\theta_{\rm ON}=0.50$ and (b) $\theta_{\rm ON}=0.90$. We employed an ERRG with $\langle k\rangle=10$ and evaluated $\rho_{\infty}$ using Eqs. (4) and (5). The red dotted line, green solid line, and blue dashed line represent the results when the adoption threshold of seed neighbors is $\theta_{\rm SN}=0.2,0.4$, and $0.6$, respectively. The symbols represent the simulation results for ERRGs with $10^{5}$ nodes, averaged over $10^{3}$ samples ($10^{2}$ graphs $\times$ $10$ initial states). The theoretical lines are in good agreement with the simulation results.
Next, we consider the case of $\theta_{\rm SN}\neq\theta_{\rm ON}$. Figures 2(a) and (b) plot the active node fraction $\rho_{\infty}$ as a function of the seed fraction $\rho_{0}$ for the extended Watts model on an ERRG with average degree $\langle k\rangle=10$, where the adoption threshold of ordinary nodes is set to $\theta_{\rm ON}=0.5$ and $0.9$, respectively. As illustrated in Fig. 2(a), when $\theta_{\rm ON}$ is low (i.e., ordinary nodes can easily become active), then when seed neighbors are activated by seeds, further activations are triggered by active seed neighbors, resulting in an activation cascade of ordinary nodes. In this case, the active node fraction jumps to $\rho_{\infty}\approx 1$ at the tipping point $\rho_{0,{\rm c}}$. In contrast, as illustrated in Fig. 2(b), when $\theta_{\rm ON}$ is high (i.e., ordinary nodes do not easily become active), activation cascades may not occur between ordinary nodes even when most seed neighbors become active. When $\theta_{\rm SN}=0.2$ and $0.4$, $\rho_{\infty}$ increases discontinuously at the tipping point but does not reach $\rho_{\infty}\approx 1$, and increases continuously as $\rho_{0}$ increases above the tipping point. At the tipping point, ordinary nodes are rarely activated, and activations spread only over the giant component formed by seed neighbors.
(a) Refer to caption (b) Refer to caption (c) Refer to caption
(d) Refer to caption (e) Refer to caption (f) Refer to caption
Figure 3: Example of how nodes of each type form a giant component at different seed fractions. The black circles, gray circles, and white circles represent seeds, seed neighbors, and ordinary nodes, respectively. (a) A network before seed placement. (b) When $\rho_{0}<\rho_{0,{\rm c}}^{\rm SN}$ (where $\rho_{0,{\rm c}}^{\rm SN}$ is the smaller critical point), only the ordinary nodes form a giant component. (c) When the seed fraction increases to $\rho_{0,{\rm c}}^{\rm SN}<\rho_{0}<\rho_{0,{\rm c}}^{\rm S}$, both seed neighbors and ordinary nodes form their own giant components. (d) When $\rho_{0,{\rm c}}^{\rm S}<\rho_{0}<\rho_{0,{\rm c}}^{\rm ON}$, seeds also form their own giant component. (e) When $\rho_{0,{\rm c}}^{\rm ON}<\rho_{0}<\rho_{0,{\rm c}}^{\rm SN}$ (where $\rho_{0,{\rm c}}^{\rm SN}$ is the larger critical point), the giant ON-component is dismantled by seeds and seed neighbors, while seeds and seed neighbors form their own giant component. (f) When $\rho_{0}>\rho_{0,{\rm c}}^{\rm SN}$, the number of seeds becomes too large, and seed neighbors also lose their giant component. As a result, only the seeds form a giant component.
@@
In the present model, it is natural to assume that the scale of cascades is related to whether nodes of each type form a giant component at different seed fractions. We refer to the connected components of seeds, the connected components of seed neighbors, the connected components of ordinary nodes, and the connected components of seed neighbors and ordinary nodes as S-components, SN-components, ON-components, and (SN+ON)-components, respectively. Whether each type of connected component forms its own giant component, which occupies a finite fraction of a network, depends on the seed fraction $\rho_{0}$ (Fig. 3). Using generating functions, we derive the size of these components and their associated critical points (Appendix A). Figure 4(a) presents the normalized giant S-component size $S_{\rm S}$, giant SN-component size $S_{\rm SN}$, giant ON-component size $S_{\rm ON}$, and giant (SN+ON)-component size $S_{\rm SN+ON}$ of the ERRG with $\langle k\rangle=10$ as a function of the seed fraction $\rho_{0}$. The giant S-component appears above a critical point $\rho_{0,{\rm c}}^{\rm S}=0.1$ when $\rho_{0}$ increases ($S_{\rm S}=0$ for $\rho_{0}\leq\rho_{0,{\rm c}}^{\rm S}$ and $S_{\rm S}>0$ for $\rho_{0}>\rho_{0,{\rm c}}^{\rm S}$). The giant ON-component and the giant (SN+ON)-component occupy most of the network when the seed fraction is small; however, their sizes decrease with $\rho_{0}$ and disappear at $\rho_{0,{\rm c}}^{\rm ON}\approx 0.207$ and $\rho_{0,{\rm c}}^{\rm SN+ON}=0.9$, respectively ($S_{\rm ON}=0$ for $\rho_{0}\geq\rho_{0,{\rm c}}^{\rm ON}$ and $S_{\rm SN+ON}=0$ for $\rho_{0}\geq\rho_{0,{\rm c}}^{\rm SN+ON}$). When $\rho_{0,{\rm c}}^{\rm ON}<\rho_{0}<\rho_{0,{\rm c}}^{\rm SN+ON}$, where $S_{\rm ON}=0$ but $S_{\rm SN+ON}>0$, the giant (SN+ON)-component is mostly composed of seed neighbors.
@@
The giant SN-component behaves non-monotonically with an increasing seed fraction $\rho_{0}$. As $\rho_{0}$ increases from zero, the number of seed neighbors first increases, and the giant SN-component is formed (Fig. 3(c)) at a smaller critical point ($\rho_{0,{\rm c}}^{\rm SN}\approx 0.011$). However, as the seed fraction continues to increase, the connectivity between seed neighbors is obstructed by seeds. As the seed fraction exceeds a larger critical point ($\rho_{0,{\rm c}}^{\rm SN}\approx 0.900$), the giant SN-component disappears (Fig. 3(f)). Thus, the giant SN-component exists only in the region where $\rho_{0}$ is between the two values of $\rho_{0,{\rm c}}^{\rm SN}$  ($S_{\rm SN}>0$ if $0.011\lessapprox\rho_{0}\lessapprox 0.900$ and $S_{\rm SN}=0$ otherwise). Figure 4(b) plots the critical point for each type of giant component in the ($\rho_{0}$, $\langle k\rangle$) plane. The figure indicates that when the average degree of the ERRG is $\langle k\rangle>\langle k\rangle_{\rm c}\approx 2.580$, each type of giant component behaves qualitatively the same as in Fig. 4(a). However, when $1<\langle k\rangle<\langle k\rangle_{\rm c}$, seed neighbors never form a giant component. In the absence of a giant SN-component, a jump in $\rho_{\infty}$ does not occur unless there is a cascade of activations between ordinary nodes. However, when the giant SN-component is present, it is possible for a cascade to occur among seed neighbors, resulting in a jump in $\rho_{\infty}$ even if ordinary nodes are rarely activated.
@@
Figure 5: (a) Active node fraction $\rho_{\infty}$ and (b) fixed points $r_{\infty}$ of Eq. (5) for the extended Watts model on an ERRG with $\langle k\rangle=10$. The adoption threshold of seed neighbors is set to $\theta_{\rm SN}=0.25$. In (a), the red dotted line, black solid line, and blue dashed line represent $\rho_{\infty}$ when $\theta_{\rm ON}=0.50$, $0.67$, and $0.90$, respectively. In (b), the red open squares, black filled circles, and blue open triangles represent $r_{\infty}$ when $\theta_{\rm ON}=0.50$, $0.67$, and $0.90$, respectively. The open circles in (a) represent the simulation results for ERRGs with $10^{5}$ nodes, averaged over $10^{3}$ samples ($10^{2}$ graphs $\times$ $10$ initial states). The simulation results are in good agreement with the theoretical line, exhibiting two jumps of $\rho_{\infty}$.
@@
Figure 6: Size distribution of active components (black squares), active SN-components (red circles), and active ON-components (blue triangles) for (a) $\rho_{0}=0.03<\rho_{0,{\rm c}}^{\rm 1st}$ (below the first tipping point), (b) $\rho_{0,{\rm c}}^{\rm 1st}<\rho_{0}=0.06<\rho_{0,{\rm c}}^{\rm 2nd}$ (above the first tipping point and below the second tipping point), and (c) $\rho_{0}=0.09>\rho_{0,{\rm c}}^{\rm 2nd}$ (above the second tipping point). The adoption thresholds of seed neighbors and ordinary nodes are set to $\theta_{\rm SN}=0.25$ and $\theta_{\rm ON}=0.67$, respectively. Each distribution, defined as the number of corresponding components of size $s$ per node, is obtained from Monte Carlo simulations of $10^{6}$ samples ($10^{4}$ initial states $\times$ $10^{2}$ graphs) for the extended Watts model on ERRGs with $10^{5}$ nodes.
We further consider the case where $\theta_{\rm SN}<\theta_{\rm ON}$, meaning that seed neighbors are more easily activated than ordinary nodes due to their proximity to seeds (the case where $\theta_{\rm SN}>\theta_{\rm ON}$ is discussed in Appendix B). The coexistence of a giant SN-component and giant ON-component suggests that a large cascade on each giant component can occur at different tipping points. Here, we demonstrate that for the extended Watts model, a tipping point exhibiting a jump in $\rho_{\infty}$ can appear twice. Let us consider the extended Watts model on an ERRG with $\langle k\rangle=10$, where a giant SN-component and giant ON-component coexist for $0.011\lessapprox\rho_{0}\lessapprox 0.207$. Figure 5(a) illustrates the active node fraction $\rho_{\infty}$ when $\theta_{\rm SN}=0.25$ and $\theta_{\rm ON}=0.50,0.67$, and $0.90$. When $\theta_{\rm ON}=0.67$ (i.e., there is a gap between the adoption thresholds of seed neighbors and ordinary nodes), there are five fixed points for $0.043\lessapprox\rho_{0}\lessapprox 0.047$, three fixed points for $0.047\lessapprox\rho_{0}\lessapprox 0.066$, and one fixed point for $\rho_{0}\gtrapprox 0.066$ (Fig. 5(b)). Consequently, the active node fraction exhibits a jump at $\rho_{0}=\rho_{0,{\rm c}}^{\rm 1st}\approx 0.047$ and at $\rho_{0}=\rho_{0,{\rm c}}^{\rm 2nd}\approx 0.066$ as $\rho_{0}$ increases.